40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Jan 4 15:11:28 2013, Jenne, Update, , TT 
    Reply  Mon Jan 14 15:15:04 2013, Jenne, Update, , TT TT2_stuck_unstuck_14Jan2013.png
       Reply  Mon Jan 14 19:56:48 2013, Manasa, Update, , TT 
          Reply  Tue Jan 15 01:41:40 2013, rana, Update, SUS, TT 
    Reply  Thu Feb 21 13:34:35 2013, Jenne, Update, Treasure, IR card removed 
Message ID: 7899     Entry time: Mon Jan 14 19:56:48 2013     In reply to: 7898     Reply to this: 7900
Author: Manasa 
Type: Update 
Subject: TT 


 [Manasa, Jenne]

The motion of the magnets (~1.5mm estimated by looking at the magnets moving) correspond to ~2deg. tilt of the mirror. This would mean almost 1.5m shift at the ETM end (~45m from the TT).

ELOG V3.1.3-